in a chemistry class 7 liters of a 4% SILVER IODIDE SOLUTION MUST BE MIXED WITH A 10% SOLUTION TO GET A 6% SOLUTION. HOW MANY LITERS OF THE 10% SOLUTION ARE NEEDED?
Answer by ptaylor(2198)

Let x=number of liters of 10% solution needed
Now we know that the pure silver iodide that exists before the mixture takes place has to equal the amount that exists after the mixture takes place, so:
0.04*7+0.10x=0.06(7+x)
0.28+0.10x=0.42+0.06x subtract 0.28 and also 0.06x from each side
0.28-0.28+0.10x-0.06x=0.42-0.28+0.06x-0.06x collect like terms
0.04x=0.14
x=3.5 liters------amount of 10% solution needed
CK
0.04*7+0.10*3.5=0.06*10.5
0.28+0.35=0.63
0.63=0.63
Hope this helps---ptaylor